Three identical small balls, each of mass 0.1g, are suspended at one point on silk thread having a length of l=20cm. What charges should be imparted to the balls for each thread to form an angle of 30 degree with the verticle.

Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
12 years ago

Dear student,

Its a  simple problem just put the values in the formula....

Mohd Ehteshaam
6 years ago
0.05x1.732x2 =a a=0.1732 meter now let the charge be q force between any two balls = kq^2 / r^2 (9x10^9)q^2/ (0.1732)^2 F=3x(10)^11(q^2) Newton now to balls will apply this force at 60* angle so vector sum of these force 2Fcos 30 (1.732x 3x(10)^11)q^2 = 5.2x(10)^11(q^2) now this force must balance component of tension in string in horizontal direction so mg tan 30* = 5.2x(10)^11(q^2) puttinng the values we get q= 0.1/3 micro coulomb or 0.03x(10)^-6